# Atom-Induced Field Squeezing Predicted by Magnus Expanding the Jaynes-Cummings Model for a Two-Level Atom

### v2

@@ Abstract
By performing up to the second order Magnus expansion to the Jaynes-Cummings Model Hamiltonian governing the interaction of a two-level atom and a single cavity mode electromagnetic field, we find the term involving the squeezing of light beyond the Rotating Wave Approximation. Our results demonstrate that the Magnus expansion provides a systematic framework to capture atom-induced squeezing effects. These findings suggest new avenues for engineering nonclassical states in simple light–matter systems without requiring multi-level structures or strong nonlinear media.
@@ I Introduction
In general, solving the Schrodinger equation involving the Hamiltonian with an explicit time-dependence is difficult. The Magnus Expansion (ME) provides an exact solution of the time evolution operator whose logarithm is expressed as a series of time-ordered integrals of nested commutators [1, 2].
To solve the time-dependent Hamiltonian governing the light-matter interaction, one usually adapts the Rotating Wave Approximation (RWA) by ignoring the contributions of high-frequency parts of either the rotated Hamiltonian or the interaction Hamiltonian [3, 4]. There are also time-averaging approaches where one finds an approximate solution for long timescales by applying a time-averaging operator on the Hamiltonian [5, 6, 7]. Such time-averaging approaches are useful for open system quantum dynamics where the quantum system is allowed to interact with complex degrees of freedom of the environment [8]. There also exists a method where one coarse-grains the Magnus evolution operator into time steps [9].
When it comes to solving the time-dependent Hamiltonian equation governing light-matter interaction between a quantized field and a two-level atom, one first rotates the time-dependent Hamiltonian into a time-independent form, using a unitary transformation that represents a rotation [10, 11]. One well-known model of quatum light-matter interaction is the Jaynes-Cummings (JC) model, which assumes that only one cavity mode of the quantum electromagnetic field interacts with the atom via weak coupling [12, 13, 14]. In this paper, we perform up to second-order Magnus Expansion on the time-evolution operator for the JC model and make comparisons to the RWA and time-averaging methods. Moreover, we demonstrate “beyond the RWA” contributions, such as the Bloch-Siegert shift and the atom-induced squeezing of light.
@@
The JC model has been a fundamental framework for quantum light-matter interaction. Phenomena beyond the RWA, such as the Bloch Siegert (BS) shift and the atom-induced field squeezing, have been studied using the JC model in the past. Various authors have analytically and numerically demonstrated field squeezing within the JC model framework [15, 16]. The BS shift, originally predicted by Bloch and Sigert in 1940 [17], has been experimentally demonstrated using circuit QED systems [18].
In this work, we analytically study the Jaynes–Cummings model without the rotating wave approximation using the Magnus expansion up to second order. While the Bloch–Siegert shift is a known feature of beyond-RWA dynamics, we find that the second-order expansion also predicts a squeezing-like interaction induced solely by a two-level atom. To the best of our knowledge, such an effect has not been derived in closed form from the JC model without additional nonlinearities or multi-level structure. This result demonstrates that field squeezing can emerge naturally from time-dependent perturbative dynamics, and offers a new analytic perspective on light–matter interactions beyond the RWA.
@@ II Methods
II.1 Transforming the Hamiltonian into the Rotating Frame
[…]
Acting $\hat{U}^{\dagger}(t)$ on both sides of the equation and using the unitary condition, one gets the following new Schrodinger equation.
[…]
Defining the “rotated” Hamiltonian as in equation 4, one gets the Schrodinger equation in the “rotated” frame as shown in equation 5.
[…]
The conditions for the solvability of general multi-level atomic systems interacting with a classical electromagnetic field under the RWA were determined in 1975 by Einwohner, Wong, and Garrison [19]. It has also been shown in our previous work that only a few of the higher-level systems are time-independent for solution via matrix exponentiation [20]. Such conditions apply to the JC model with general multi-level atomic systems, as rotating the field contributes in eliminating the field term and inducing frequency-dependent terms, as we will see in section II.3. The analysis for the JC model for general multilevel atomic systems is beyond the scope of this research.
II.2 The Magnus Expansion
In general, solving a system with a time-dependent Hamiltonian is a challenging task. One approach is to rotate the Hamiltonian into a frame where the Hamiltonian is time-independent and solve the Hamiltonian under the RWA. Even under the RWA, only a few systems can be rotated into a time-independent form [20]. However, there exists a method that gives exact solutions to any time-dependent Schrodinger equation involving a Hamiltonian with explicit time-dependence on the condition that the expansion converges [21]. Such an expansion is called the Magnus Expansion.
@@ II.2 The Magnus Expansion
II.3 The Jaynes-Cummings Model in the Rotating Frame
@@
In the above equation, $\hat{a}^{\dagger}$ and $\hat{a}$ represents the creation and annihilation operators of the cavity mode of interest, $\omega$ and $\omega_{0}$ the frequency of the mode and that of the atom, $\hat{\sigma}_{z}$ the Pauli-z operator for a two-level atom, $\hat{\sigma}_{\pm}$ the atomic raising and lowering operators, and g the coupling constant between the electromagnetic field and the atom transition. The matrix representation of the $\hat{\sigma}$ operators is as follows.
$\hat{\sigma}_{z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},\hat{\sigma}_{-}=\begin{pmatrix}0&0\\ 1&0\end{pmatrix},\hat{\sigma}_{+}=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}$ (10)
In standard calculations of the JC model, we perform two rotations: one for the field and one for the atom. We first rotate the atom part of the Hamiltonian by using the unitary matrix $\hat{U}(t)=\exp(-i\omega_{0}\hat{\sigma}_{z}t/2)$. Notice that $i\hbar\hat{U}^{\dagger}(t)\frac{d\hat{U}(t)}{dt}=-\hbar\omega_{0}/2\hat{\sigma}_{z}$ cancels with $\hat{H}_{at}(t)$. The Pauli-z operator $\hat{\sigma}_{z}$ commutes with the creation, annihilation operators, and itself. Using the unitary condition, we get $\hat{U}^{\dagger}(t)(\hat{H}_{F}(t)+\hat{H}_{at}(t))\hat{U}(t)=\hat{H}_{F}(t)+\hat{H}_{at}(t)$. The interaction part of the Hamiltonian gets rotated as $\hat{U}^{\dagger}(t)\hat{H}_{I}(t)\hat{U}(t)=\hbar g(\hat{a}^{\dagger}-\hat{a})\hat{\sigma}_{-}e^{-i\omega_{0}t}+(\hat{a}^{\dagger}-\hat{a})\hat{\sigma}_{+}e^{i\omega_{0}t}$. Hence, the Hamiltonian after the atomic rotation is determined as follows.
$\hat{H}_{rot}(t)=\hbar\omega\hat{a}^{\dagger}\hat{a}+i\hbar g(\hat{a}^{\dagger}-\hat{a})(\hat{\sigma}_{-}e^{-i\omega_{0}t}+\hat{\sigma}_{+}e^{i\omega_{0}t})$ (11)
Next, we do a second rotation by introducing the field-rotation unitary $\hat{V}(t)=\exp(-i\omega\hat{n}t)$ where $\hat{n}=\hat{a}^{\dagger}\hat{a}$ is the hermitian photon number operator. Due to the hermiticity, $\hat{V}^{\dagger}(t)=\exp(i\omega\hat{n}t)$. We get $i\hbar\hat{V}^{\dagger}(t)\frac{d\hat{V}(t)}{dt}=-\hbar\omega\hat{n}$ and hence, the first term in $\hat{H}_{rot}(t)$ is cancelled. Using the Baker-Campbell-Haussdorff (BCH) lemma (see appendix A), we get the relation $\hat{V}^{\dagger}(t)\hat{a}\hat{V}(t)=\hat{a}e^{-i\omega t}$ and its conjugate $\hat{V}^{\dagger}(t)\hat{a}^{\dagger}\hat{V}(t)=\hat{a}^{\dagger}e^{i\omega t}$. Combining everything, we get the fully-rotated Hamiltonian as follows.
@@
Here, we defined detuning to be $\Delta=\omega-\omega_{0}$ and the sum-frequency $\Sigma=\omega+\omega_{0}$. Notice that the “rotated” Hamiltonian is similar to the interaction picture Hamiltonian. However, we are staying in the full Schrödinger picture, as the operators hold no time dependence.
Notice the terms oscillating at frequency $\Sigma$. Such terms do not contribute much to system dynamics. Physically, the term oscillating at $\Sigma$ destroys a photon and lowers the atom from excited to ground state, while the term oscillating at $\Sigma$ creates a photon and excites the atom. These processes do not obey the conservation of energy. Nevertheless, under short timescales, the energy-time uncertainty relations allow rapid fluctuations in system dynamics. Many theoretical derivations and experimental applications have neglected such highly oscillating terms by making a so-called “Rotating Wave Approximation (RWA)” [24]. Moreover, it has also been shown that one must be careful when using the RWA as the approximation can produce non-causal behaviors [25, 26, 27, 28]. There also exists a paper by Bugarth and colleagues on the effects of the RWA and the experimental validity of the approximation [29]. In this work, we briefly go over the “averaging” process of the first-order Magnus expansion and its relation to the RWA. We also capture non-trivial contributions from the terms neglected by the RWA in section III.2.
@@ III.2 The Second Order Magnus Expansion
We have to find the commutator at different times. First, we write down the well-known commutation relations.
$\begin{split}&[\hat{a},\hat{a}^{\dagger}]=1,\\ &[\hat{\sigma}_{-},\hat{\sigma}_{+}]=-\hat{\sigma}_{z}\end{split}$ (16)
Noting that every commutator commutes with itself, we get the commutator at two different times as follows.
$\begin{split}-\frac{[\hat{H}^{\prime}(t_{1}),\hat{H}^{\prime}(t_{2})]}{\hbar^{2}g^{2}}=&[\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}\hat{\sigma}_{+}](-e^{i\Delta(t_{1}-t_{2})}+e^{-i\Delta(t_{1}-t_{2})})+\\ &[\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}^{\dagger}\hat{\sigma}_{+}](e^{i(\Delta t_{1}+\Sigma t_{2})}-e^{i(\Sigma t_{1}+\Delta t_{2})})+\\ &[\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}\hat{\sigma}_{-}](-e^{i(\Delta t_{1}-\Sigma t_{2})}+e^{-i(\Sigma t_{1}-\Delta t_{2})})+\\ &[\hat{a}^{\dagger}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{+}](-e^{i(\Sigma t_{1}-\Delta t_{2})}+e^{-i(\Delta t_{1}-\Sigma t_{2})})+\\ &[\hat{a}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{-}](e^{-i(\Delta t_{1}+\Sigma t_{2})}-e^{-i(\Sigma t_{1}+\Delta t_{2})})+\\ &[\hat{a}^{\dagger}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{-}](-e^{i\Sigma(t_{1}-t_{2})}+e^{-i\Sigma(t_{1}-t_{2})})\end{split}$ (17)
Since the commutators do not have explicit time dependence, we do the time integrals first and see the contributions of each term present in equation 17. The time integrals are as follows.
$I_{1}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(-e^{i\Delta(t_{1}-t_{2})}+e^{-i\Delta(t_{1}-t_{2})})$
@@
$I_{2}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(e^{i(\Delta t_{1}+\Sigma t_{2})}-e^{i(\Sigma t_{1}+\Delta t_{2})})$
@@
$I_{3}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(-e^{i(\Delta t_{1}-\Sigma t_{2})}+e^{-i(\Sigma t_{1}-\Delta t_{2})})$ (20)
$I_{4}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(-e^{i(\Sigma t_{1}-\Delta t_{2})}+e^{-i(\Delta t_{1}-\Sigma t_{2})})$ (21)
$I_{5}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(e^{-i(\Delta t_{1}+\Sigma t_{2})}-e^{-i(\Sigma t_{1}+\Delta t_{2})})$
$=\frac{\omega_{0}e^{-2i\omega t}-\omega e^{-i\Sigma t}+\omega e^{-i\Delta t}-2\omega_{0}}{(\omega^{2}-\omega_{0}^{2})\omega}=I_{2}^{*}$ (22)
$I_{6}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(-e^{i\Sigma(t_{1}-t_{2})}+e^{-i\Sigma(t_{1}-t_{2})})$
@@
Note that $I_{3}$ and $I_{4}$ are not evaluated because the commutators associated with these integrals are zero. Also, $I_{2}$ and $I_{5}$ being complex conjugates of each other is crucial in yielding the atomic squeezing term later. For notation, call $I_{2}=\zeta$ and $I_{5}=\zeta^{*}$.
@@
$[\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}\hat{\sigma}_{-}]=-\hat{\sigma}_{-}^{2}=0$ (27)
$[\hat{a}^{\dagger}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{+}]=-\hat{\sigma}_{+}^{2}=0$ (28)
@@
Notice that the second-order Magnus operator is manifestly anti-hermitian as the first two terms involve the product of a purely imaginary number and a hermitian operator, while the last term is the well known anti-hermitian squeezing operator. This agrees with the condition that every order of $\hat{\Omega}(t)$ must be anti-hermitian to maintain unitarity of the Magnus time evolution operator. The term $i\frac{g^{2}t}{\Delta}\hat{a}^{\dagger}\hat{a}\hat{\sigma}_{z}$ represents a process where the atomic energy is shifted according to the number of photons present. This can be interpreted as an AC-Stark Shift. Since $|\hat{\sigma}_{+}\hat{\sigma}_{-}=\ket{1}\bra{1}$, it represents the frequency shift of the excited state. The term $i\frac{g^{2}t}{\Sigma}\hat{a}^{\dagger}\hat{a}\hat{\sigma}_{z}$ represents the Bloch-Siegert shift as it is a small photon-induced energy shift suppressed by a large frequency $\omega+\omega_{0}$. Moreover, our results agree with Wall’s prediction that the Bloch-Siegert shift scales with the number of photons, which was later verified experimentally [31, 32].
The last term involves the $\zeta^{*}\hat{a}^{2}-\zeta(\hat{a}^{\dagger})^{2}$ term, which resembles the squeezing operator of the quantized electromagnetic field with $\zeta=\frac{\omega_{0}e^{2i\omega t}-\omega e^{i\Sigma t}+\omega e^{i\Delta t}-\omega_{0}}{\omega(\omega^{2}-\omega_{0}^{2})}$ [14]. Therefore, it can be interpreted as the presence of the atom induces some squeezing towards the existing light field. This behavior would not be captured with RWA, as the cross-contribution of rotating and counterrotating terms is responsible for squeezing. Moreover, the amplitude of the squeezing term is not negligible when the laser is tuned near resonance, and it seems like it diverges on resonance. However, the term does not diverge on resonance, and the proof is shown in appendix B. Inspecting the limit shown in appendix B, it can also be noted that the amplitude is quite small when the system is tuned near resonance. One interesting remark is that the coefficient of the squeezing term is an exponent that is on another exponent due to the nature of the Magnus evolution operator. This means that the squeezing effect fluctuates very rapidly in time. Therefore, it is interesting to inspect the contribution of such a term in atom dynamics for potential experimental evidence.
Notice the increase in powers of the coupling constant in the Magnus expansion. The Magnus expansion converges when $gt<<\pi$ [2]. The standard values for coupling constants for cavity QED setups are $g=10^{6}-10^{7}$Hz [29]. For long timescale quantum dynamics, coarse-graining the Magnus expansion on periods on the orders below the microsecond scale has been suggested [9].

### v4

# Conditional squeezing induced by a two-level system: a first-principles approach to QND readout

@@ Abstract
We present a systematic Magnus expansion treatment of light-matter interaction beyond the Rotating Wave Approximation. We show that at the second order of Magnus series, the time-evolution operator acquires both energy-shifts and squeezing contributions. In addition to the energy shifts caused by vacuum and photon numbers, the second-order evolution operator contains a term that induces conditional squeezing of the field mode depending on the state of the atom. Such a term suggests a natural mechanism for phase-sensitive, quantum non-demolishing type measurements of squeezed light via homodyne detection. We also show that the second-order Magnus operator in a close SU(1,1) algebra, ensuring the exponentiation of the Magnus series yields a well-defined unitary evolution. By deriving squeezing directly from the Jaynes-Cummings Hamiltonian, our results clarify how energy shifts and $\hat{\sigma}_{z}$-dependent squeezing arise from an SU(1,1) structure, with direct implications for phase-sensitive and quantum nondemolition measurements.
Magnus expansion, non-demolishing measurements, quadrature squeezing, Jaynes-Cummings model, beyond the Rotating Wave Approximation (RWA), SU(1,1) group
@@ I Introduction
Squeezed states of light, first predicted theoretically to reduce quantum fluctuations beyond the shot-noise limit [6], were later experimentally realized by four-wave mixing in an optical cavity [23]. Such states were later incorporated in enhanced precision measurements in interferometric setups [28, 25, 22]. Beyond laboratory-scale demonstrations, squeezed states have also been applied in large-scale interferometric detectors. The LIGO collaboration has demonstrated the use of squeezed light to enhance the sensitivity of gravitational detectors [25].
Most measurements in quantum information destroy the state of the qubit due to the collapse postulate. It is important to have a way of performing measurements from a quantum system without irreversibly disturbing the observable of interest. Such measurements are called quantum nondemolishing (QND) measurements [19, 4]. In this article, we demonstrate how the interaction of a simple two-level system with a single-mode field opens possibilities of performing such measurements.
@@
We perform up to second-order Magnus Expansion on the time-evolution operator for the Jaynes-Cummings(JC) model [11]. We demonstrate “beyond the Rotating Wave Approximation(RWA)” contributions, such as the Bloch-Siegert shift and the atom-induced conditional squeezing of light. The JC model has been a fundamental framework for quantum light-matter interaction. Various authors have analytically and numerically demonstrated field squeezing within the JC model framework [12, 30]. The Bloch-Siegert shift, originally predicted by Bloch and Siegert in 1940 [3], has been experimentally demonstrated using circuit QED systems [8]. While the Bloch–Siegert shift is a known feature of beyond-RWA dynamics, we find that the second-order expansion also predicts a two-photon term governing the squeezing of light induced solely by a two-level atom.
Although squeezing effects in two-level systems have been discussed in the literature, our approach derives the effect in closed form from first principles, without recourse to multilevel extensions or external nonlinearities. Traditionally, squeezed light is generated in nonlinear optical media using techniques such as degenerate parametric down-conversion in $\chi^{(2)}$ crystals in an optical cavity [29, 31] and four-wave mixing [24, 21]. By contrast, we demonstrate that even the simplest Jaynes–Cummings two-level atom intrinsically contains a squeezing term when treated beyond the RWA. While the magnitude of the squeezing is not expected to compete with engineered nonlinear media, the finding highlights an overlooked resource. In particular, deriving a squeezing coefficient directly from first principles strengthens the value of theoretical quantum optics, where simplistic models not only predict nonclassical light but also provide a foundation for understanding how such effects can be systematically extended to more complex, scalable systems.
This paper is arranged as follows. In Section II, we present a brief discussion of textbook materials such as the Magnus Expansion and the Jaynes-Cummings Model. In Section III, we present our calculation of the second-order Magnus operator and its implications toward QND Measurement by homodyne detection of squeezed light. Finally, Section IV summarizes our conclusions and highlights future work.
@@ II Methods
II.1 The Magnus Expansion
@@
II.2 The Jaynes-Cummings Model in the Rotating Frame
@@
Here $\hbar$ represents the Planck’s constant divided by $2\pi$, i the imaginary number, g the coupling strength between the atom and the field, $\omega$ the frequency of the cavity field mode, $\omega_{0}$ the transition frequency of the two-level system, $\hat{a}^{\dagger}$ and $\hat{a}$ the photon creation and annihilation operators f the field, $\hat{\sigma}_{z}$ the Pauli-z operator of the atom, and $\hat{\sigma}_{+}$ and $\hat{\sigma}_{-}$ the raising and lowering operator in the atomic Hilbert space. The term $h.a.$ represents the Hermitian adjoint of the operator in brackets. Defining the “rotated” Hamiltonian as in equation 5, one can write down the Schrodinger equation in the “rotated frame” [9].
Note that we stay in Schrodinger’s picture in this article as only the state gains time-dependence via a unitary rotation. The operators obtain no time dependence in contrast to the interaction picutre.
In standard calculations of the JC model, we perform two rotations: one for the field and one for the atom [9]. The unitary operator representing the atomic rotation is $\hat{U}(t)=\exp(-i\omega_{0}\hat{\sigma}_{z}t/2)$ and that of the field is $\hat{V}(t)=\exp(i\omega\hat{n}t)$ where $\hat{n}$ is the hermitian photon number operator defined as $\hat{n}=\hat{a}^{\dagger}\hat{a}$. After the Hamiltonian rotations using equation 5, we get the fully rotated Jaynes-Cummings Hamiltonian as follows.
@@
Here, we defined detuning to be $\Delta=\omega-\omega_{0}$ and the sum-frequency $\Sigma=\omega+\omega_{0}$. Notice the terms oscillating at frequency $\Sigma$. Such terms do not contribute much to system dynamics. Physically, the term oscillating at $\Sigma$ destroys a photon and lowers the atom from excited to ground state, while the term oscillating at $\Sigma$ creates a photon and excites the atom. These processes do not obey the conservation of energy. Nevertheless, under short timescales, the energy-time uncertainty relations allow rapid fluctuations in system dynamics. Many theoretical derivations and experimental applications have neglected such highly oscillating terms by making a so-called “Rotating Wave Approximation (RWA)” [18]. Time-averaging effective Hamiltonian theories have also made similar approximations to the RWA by neglecting higher-order fluctuations by taking ensemble averages of the dynamics [10]. Moreover, it has also been shown that one must be careful when using the RWA as the approximation can produce non-causal behaviors [16, 15, 20, 7]. There also exists a paper by Bugarth and colleagues on the effects of the RWA and the experimental validity of the approximation [5]. In this work, we briefly go over the “averaging” process of the first-order Magnus expansion and its relation to the RWA. We also capture non-trivial contributions from the terms neglected by the RWA in section III.2.
@@ III.2 The Second Order Magnus Expansion
We have to find the commutator at different times. Noting that every commutator commutes with itself, we get the commutator at two different times as follows.
$\begin{split}-\frac{[\hat{H}^{\prime}(t_{1}),\hat{H}^{\prime}(t_{2})]}{\hbar^{2}g^{2}}&=[\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}\hat{\sigma}_{+}](-e^{i\Delta(t_{1}-t_{2})}+e^{-i\Delta(t_{1}-t_{2})})+\\ &[\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}^{\dagger}\hat{\sigma}_{+}](e^{i(\Delta t_{1}+\Sigma t_{2})}-e^{i(\Sigma t_{1}+\Delta t_{2})})+\\ &[\hat{a}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{-}](e^{-i(\Delta t_{1}+\Sigma t_{2})}-e^{-i(\Sigma t_{1}+\Delta t_{2})})+\\ &[\hat{a}^{\dagger}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{-}](-e^{i\Sigma(t_{1}-t_{2})}+e^{-i\Sigma(t_{1}-t_{2})})\end{split}$ (10)
Here, we neglect terms involving $\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}\hat{\sigma}_{-}$] and $\hat{a}^{\dagger}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{+}$] as the commutators evaluate to zero. Since the commutators do not have explicit time dependence, we do the time integrals first and see the contributions of each term present in equation 10. The time integrals are as follows.
$I_{1}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(-e^{i\Delta(t_{1}-t_{2})}+e^{-i\Delta(t_{1}-t_{2})})$ (11)
@@
$I_{2}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(e^{i(\Delta t_{1}+\Sigma t_{2})}-e^{i(\Sigma t_{1}+\Delta t_{2})})$ (13)
@@
$I_{3}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(e^{-i(\Delta t_{1}+\Sigma t_{2})}-e^{-i(\Sigma t_{1}+\Delta t_{2})})$ (15)
$=\frac{\omega_{0}e^{-2i\omega t}-\omega e^{-i\Sigma t}+\omega e^{-i\Delta t}-\omega_{0}}{(\omega^{2}-\omega_{0}^{2})\omega}=I_{2}^{*}$ (16)
$I_{4}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(-e^{i\Sigma(t_{1}-t_{2})}+e^{-i\Sigma(t_{1}-t_{2})})$ (17)
@@
Note that $I_{2}$ and $I_{3}$ being complex conjugates of each other is crucial in yielding the atomic squeezing term later. For notation, call $I_{2}=\zeta(t)/g^{2}$ and $I_{3}=\zeta^{*}(t)/g^{2}$.
@@
Notice that the second-order Magnus operator is manifestly anti-hermitian as the first two terms involve the product of a purely imaginary number and a hermitian operator, while the last term is the well known anti-hermitian squeezing operator formally defined as $\exp(\frac{1}{2}(\zeta^{*}(t)\hat{a}^{2}-\zeta(t)(\hat{a}^{\dagger})^{2})$. This agrees with the condition that every order of $\hat{\Omega}(t)$ must be anti-hermitian to maintain unitarity of the Magnus time evolution operator.
In equation 24, we can rewrite $|\ket{1}\bra{1}$ as $\frac{1}{2}(\hat{I}_{at}+\hat{\sigma}_{z})$ and $|\ket{0}\bra{0}$ as $\frac{1}{2}(\hat{I}_{at}-\hat{\sigma}_{z})$ where $\hat{I}_{at}$ is the identity operator in the Hilbert Space of the two-level atom. Noting that the identity operator commutes with every other operator, we can ignore the contribution of the term associated with the identity operator as it only adds a global phase to the system. Hence, we rewrite equation 24 as follows.
$\begin{split}\hat{\Omega}_{2}(t)=&ig^{2}(\hat{a}^{\dagger}\hat{a}+\frac{1}{2})(\frac{t}{\Delta}-\frac{t}{\Sigma}+\frac{\sin(\Sigma t)}{\Sigma^{2}}-\frac{\sin(\Delta t)}{\Delta^{2}})\hat{\sigma}_{z}\\ &+\frac{1}{2}\left(\zeta^{*}(t)\hat{a}^{2}-\zeta(t)(\hat{a}^{\dagger})^{2}\right)\hat{\sigma}_{z}\end{split}$ (25)
The first term represents the time-dependent frequency shifts, involving the AC-stark shift governed by terms involving $\Delta$, the Bloch-Siegert shift governed by terms involving $\Sigma$. The last term involves the $\frac{1}{2}(\zeta^{*}\hat{a}^{2}-\zeta(\hat{a}^{\dagger})^{2})$ term, which resembles the squeezing operator of the quantized electromagnetic field with the closed-form expression of the squeezing coefficient presented in equation 26 [9].
$\zeta(t)=g^{2}\frac{\omega_{0}e^{2i\omega t}-\omega e^{i\Sigma t}+\omega e^{i\Delta t}-\omega_{0}}{\omega(\omega^{2}-\omega_{0}^{2})}$ (26)
Therefore, it can be interpreted as the presence of the atom induces some squeezing towards the existing light field. This behavior would not be captured with RWA, as the cross-contribution of rotating and counterrotating terms is responsible for squeezing. We note that exponentiation of Equation 25 requires care due to non-commuting operators; details are given in Appendix A.
Moreover, the dependence of the $\hat{\sigma}_{z}$ operator on the squeezing term means that the squeezed light is going to have a $\pi$ phase shift depending on whether the atom is on the ground state or the excited state. There exists a possibility that one can measure the light interacting with a two-level system in a homodyne detector to possibly determine the state of the atom using quantum non-demolition (QND) measurements for quantum information [19, 4].